OpenStudy (anonymous):

Multiply: z squared plus 6 times z plus 5 all over 3. • 9 times x all over the quantity z squared plus 5 times z, end quantity.

OpenStudy (anonymous):

multiple choice, 3 times the quantity z plus 2 all over z., z ≠ 0, −5 3 times x times the quantity z plus 1 all over z., z ≠ 0, −5 6 times z plus 3 times x all over z., z ≠ −5 18 times x times z all over the quantity 3 plus z, end quantity., z ≠ −5

OpenStudy (ipwnbunnies):

\[\frac{z^2 + 6z + 5}{3} * \frac{9x}{z^2 + 5z}\]

OpenStudy (anonymous):

yep, thats it

OpenStudy (ipwnbunnies):

If you see, the first step is the get that '3' out the denominator. You can get it out by dividing the 9x by 3, which will give you 3x. It's the same as dividing the whole numerator by 3. \[\frac{9(z^2 + 6z + 5)}{z^2 + 5z}\]

OpenStudy (anonymous):

next step?

OpenStudy (ipwnbunnies):

Since anything divided by 0 is undefined, x(x+5) cannot equal 0. If we solve for x, that means x = 0 and -5. The function is not defined at x = 0 and -5.

OpenStudy (ipwnbunnies):

\[\frac{3x(z+1)(z+5)}{z(z+5)}\]

OpenStudy (ipwnbunnies):

I meant z cannot equal 0 or -5. Argh. You can see that a (z+5) factor can cancel out \[\frac{3x(z+1)}{z}, z \neq 0, -5\]

OpenStudy (anonymous):

yes